OpenStudy (anonymous):

write the summation without sigma notation. then evaluate them. will be drawn inside! @hartnn @larryboxaplenty did i do it right? :/

OpenStudy (anonymous):



OpenStudy (anonymous):

so would i get this? 1-1/1 = 0 2-1/2 = 1/2 3-1/2 = 2/2 = 1 ? so 0 + 1/2 + 1 ?

OpenStudy (anonymous):

and then you evaluate and get 1.5 ?

OpenStudy (anonymous):

You should get 1.16, with the 6 repeating = 7/6

OpenStudy (anonymous):

oh :/ darn so i did it wrong hahaa

OpenStudy (anonymous):

(1-1)/1 = 0/1 = 0 (2-1)/2 = 1/2 = 0.5 (3-1)/3 = 2/3 0+0.5+(2/3) = 1.16 repeating = 7/6

OpenStudy (anonymous):

You were right, until the 3rd term. You divided by 2 instead of 3

OpenStudy (anonymous):

ahh okay i see :P so because you get basically 3/6 + 4/6 = 7/6 so 7/6 would be the final answer? and it would be 0+ 1/2 + 2/3 ?

OpenStudy (anonymous):

Yes, (1/2) = 3/6 and (2/3) = 4/6

OpenStudy (anonymous):

ahh okay awesome! thank you! :D